OpenStudy (anonymous):

to solve for a variable involving an nth root, raise both sides of the equation to the nth power (y√n)=y f(x)=x√+2 3 minutes agoReport Abuse dessyk 0 (y√n)n=y x√+2

OpenStudy (turingtest):

you gotta use the equation feature, that question is incoherent.

OpenStudy (anonymous):

\[(\sqrt[n]{y})^n =y\] problem is; \[f(x)= \sqrt{x} +2\]

OpenStudy (anonymous):